# Synthesizing Sheet Music Problems for Evaluation and Reinforcement Learning

### v1

@@ 4.3 RLVR on Synthetic Music Reasoning Data
Compared to previous approaches, our data synthesis framework provides a simple yet effective method for scaling sheet music reasoning data with verifiable answers, enabling RLVR on LLMs and MLLMs. We train two distinct models on our synthetic dataset of 8,000 examples. The text-based Qwen3-8B-Base is trained on textual data, while the vision-language model, Qwen2.5-VL-7B-Instruct, is trained on visual data. Both models are optimized using the GRPO algorithm. Table 3 presents the performance of these models on our SSMR-Bench.
In the Textual Modality, the Qwen3-8B-Base model initially achieved relatively low scores across all metrics, with an average score of 23.18%, prior to the application of GRPO. However, after incorporating GRPO on the synthetic training set, performance shows significant improvements, with average scores increasing by 47.76% points, resulting in an impressive final score of 70.94%. This also represents a 13.06% improvement over the baseline Qwen3-8B score of 57.88%. In the Visual Modality, the Qwen2.5-VL-7B-Instruct model initially recorded moderate results with an average score of 41.81%, prior to applying RLVR. Following RL training, the model’s performance improved by an average of 28.75% points, reaching a final score of 70.56%.
Table 3: Performance of Qwen3-8B-Base and Qwen2.5-VL-7B-Instruct on the SSMR-bench after applying GRPO on synthetic data, with results compared to the Qwen3-8B and Qwen2.5-VL-7B-Instruct, respectively. The table also highlights the improvements.
Models Synthesize Sheet Music Reasoning Benchmark
@@
4.4 Results Analysis by Question Type
Figure 4: Performance comparison of various selected LLMs and MLLMs on different music-related question types. The upper plot shows the performance of LLMs evaluated on the textual modality, while the lower plot displays the performance of MLLMs in the visual modality. The performance is assessed across nine distinct question categories, such as Bar Placement, Chord Identification (Chord ID), and Scale Identification (Scale ID). Detailed descriptions of these question types can be found in Appendix A.
The upper plot of Figure 4 illustrates the performance of LLMs in the textual modality. After reinforcement learning on 8,000 synthetic music reasoning samples through GRPO, Qwen3-8B-Base outperforms Qwen3-8B in 2/3 of the question types, despite the latter undergoing extensive data training. Additionally, Qwen3-8B-Base+GRPO achieves comparable performance to Deepseek-V3 and Qwen3-32B in aspects such as "Time Sig", "Interval No", and "Note Comp". However, for "Chord ID" and "Scale Sel," the model shows limited improvement, which can be attributed to the scarcity of these data types in the training set shown in Appendix A.
5 Beyond Synthetic Data and Question Answering Tasks
5.1 RL on Synthetic Data Benefits Real-World Scenarios
To assess the effectiveness of our synthetic data, we conduct experiments on MusicTheoryBench (Yuan et al., 2024). Specifically, we evaluate the performance of Qwen3-8B-Base after GRPO on the synthetic sheet music problems and compare it against three baselines: the unmodified Qwen3-8B-Base, the officially released reasoning model Qwen3-8B, and a GRPO-trained variant using MATH data. Additionally, we report results for GPT-4, GPT-4 with Role Play, and GPT-4 with Chain-of-Thought (CoT) reasoning (Yuan et al., 2024) to provide further context. The results are summarized in Table 4. Besides, we also evaluate the Qwen3-8B-Base model and its variants on a series of math benchmarks, and present results in Table 5.
As shown in Table 4, Qwen3-8B-Base demonstrates substantial performance gains after GRPO training on the synthetic data. Its original performance was 37.92% in knowledge and 19.39% in reasoning, averaging 28.65%. After GRPO, the model improves by 25.28% in knowledge and 17.34% in reasoning, corresponding to an average gain of 21.32%. Compared to Qwen3-8B, the improvement is mainly observed in reasoning (up by 7.14%) with a marginal increase in knowledge (0.37%). In contrast, the Qwen3-8B-Base variant trained with GRPO on MATH data exhibits much smaller gains, achieving only a 33.81% average score, which is 12.4% lower than Qwen3-8B. Besides, Qwen3-8B-Base reinforced with our synthetic data outperforms GPT-4 in overall performance and achieves reasoning capabilities comparable to GPT-4 with Role Play and GPT-4 with CoT.
This confirms the effectiveness of synthesizing verifiable questions specifically related to sheet music in enhancing the model’s understanding and reasoning abilities within this domain, while yielding only moderate gains in Math problem-solving. In contrast, applying GRPO to MATH data yields substantially smaller gains in sheet music reasoning while producing significant improvements in mathematical tasks, underscoring the critical importance of domain relevance in synthetic data generation. Overall, these findings suggest that automatically generating verifiable sheet music problems provides a practical and effective approach for strengthening reasoning capabilities in sheet music comprehension.
Table 4: Evaluation results on the MusicTheoryBench. Training the Qwen3-8B-Base model with GRPO and our synthetic data leads to substantial gains in both knowledge and reasoning. Notably, our enhanced model also outperforms the Qwen3-8B, especially in the reasoning category.
Models MusicTheoryBench
Knowledge Reasoning Avg.
@@ 5.1 RL on Synthetic Data Benefits Real-World Scenarios
Qwen3-8B $62.83$ $29.59$ $46.21$
@@
   + Math GRPO 47.2147.21(+9.29) 20.0420.04(+0.65) 33.8133.81(+5.16)
   + Music GRPO 63.2063.20(+25.28) 36.7336.73(+17.34) 49.9749.97(+21.32)
Models Math Benchmarks
AIME24 AMC MATH-500 Minerva Olympiad Bench
Qwen3-8B $50.41$ $71.08$ $91.60$ $45.96$ $60.89$
Qwen3-8B-Base $8.75$ $39.16$ $66.00$ $25.37$ $30.22$
+ Math GRPO 25.8325.83(+17.08) 62.0462.04(+22.88) 85.4085.40(+19.40) 40.4440.44(+15.07) 50.5250.52(+20.30)
+ Music GRPO 13.7513.75(+5.00) 52.1152.11(+12.95) 79.4079.40(+13.40) 34.1934.19(+8.82) 40.4440.44(+10.22)
5.2 From Question Answering to Sheet Music Generation
@@
We introduce two primary evaluation metrics:
  • Rhythmic Consistency (RC): The average of a binary score assigned to each generated continuation. A sample is scored 1 if all four measures are rhythmically correct and 0 otherwise.
  • Novel Rhythmic Coherence (NRC): While melody repetition is common in music, we also aim to assess a model’s likelihood of producing rhythmically correct sheet music when generating novel melodies. The NRC metric is designed for this purpose. It is calculated as the RC score, but only for generated continuations that are not direct copies of the provided prompt.
Figure 5: Comparison of rhythmic consistency. The figure shows the Rhythmic Consistency (RC) and Novel Rhythmic Coherence (NRC) for Qwen3-8B-Base and its GRPO variant, trained on our synthetic data.
@@
We randomly sample 30 pieces of sheet music from SSMR-bench and evaluate each model’s continuation through human judgment. The results in Table 5 show that training on synthetic sheet music problems also leads to improvements in music composition. The base Qwen3-8B model achieves an RC score of 66.67%. After reinforcement learning on our synthetic dataset, this score dramatically increases to 99.33%, surpassing the 70.00% from the original Qwen3-8B. The improvement in generating novel content is also evident: our fine-tuned model’s NRC score of 88.89% is a significant increase over the base model’s 50.00% and the original Qwen3-8B’s 68.97%. Although we focus only on the rhythm dimension, which is just one aspect of music, the observed improvements also highlight the potential for enhancing reasoning abilities to facilitate music composition.
6 Discussion
Possible Challenges Facing AI Musicians.
The findings presented in Section 4.2 highlight the varying levels of difficulty associated with different modalities of sheet music. In particular, the same question, when presented in visual sheet music format, is significantly more challenging for LLMs than when conveyed in a textual format. Beyond modality, sheet music can be represented in a variety of formats, including MIDI, MusicXML, and others. Furthermore, sheet music is not limited to traditional staff notation; alternative formats, such as Guitar Tablature and numbered musical notation, should also be considered, which together present substantial challenges for the development of an ideal AI musician.
Limitations and Future Work.
In this work, we focus on synthesizing verifiable sheet music reasoning problems based on widely accepted rules in music theory. We implement this approach through a data generation framework and evaluate its effectiveness. However, the questions we construct are relatively simple, and can be reliably solved by Deepseek-R1. At the same time, they highlight the challenge for MLLMs to fully understand music sheets, as evidenced by the low performance of Gemini2.5-Pro. Taken together, these observations motivate future work on designing more complex, rule-based, verifiable music reasoning questions and on developing improved methods to enhance models’ reasoning abilities of sheet music, as well as transferring these abilities to music composition.
@@ 7 Conclusion
In this work, we introduce the concept of synthesizing sheet music problems based on the rules of music theory, enabling the construction of verifiable questions for evaluation and training. We implement this idea by proposing a framework that automatically generates sheet music questions with staff notation in both textual and visual modalities. Based on this framework, we introduce SSMR-Bench and a corresponding training set in both modalities. Experimental results demonstrate that the synthesized data effectively enhances the reasoning capabilities of LLMs and MLLMs in sheet music tasks, underscoring its potential as a valuable resource. Moreover, the results also highlight the promise of this enhanced ability to facilitate music composition. Our work demonstrates the potential of programmatically generating verifiable sheet music problems to evaluate and improve the abilities of LLMs and MLLMs in sheet music, while also highlighting promising avenues for future research toward AI musicianship.
@@ Appendix A Dataset Details
A.1 Dataset Distribution
@@
A.2 Question Class Descriptions
@@ A.2.1 Domain: Scale
A.2.2 Domain: Rhythm
This domain focuses on rhythm, time signatures, and measure structure.
@@ A.2.2 Domain: Rhythm
A.2.3 Domain: Interval
This domain tests the ability to identify and construct intervals between notes.

### v2

@@ 3.3 RLVR on Synthetic Music Reasoning Data
Table 3: Performance of Qwen3-4B-Base, Qwen3-8B-Base, and Llama3.1-8B-IT on Textual QA, and Qwen2.5-VL-3B-IT and Qwen2.5-VL-7B-IT on Visual QA tasks, after applying GRPO to synthetic data on the SSMR-bench. The table compares these results with their baseline models (without GRPO) and the corresponding Thinking models (if applicable).
Models Synthetic Sheet Music Reasoning Benchmark
@@
Compared to previous approaches, our data synthesis framework offers a simple yet effective method for scaling sheet music reasoning data with verifiable answers, enabling RLVR on both LLMs and MLLMs. We select various open-source LLMs and VLMs and train them on our synthetic dataset consisting of 8,000 examples. The LLMs, such as Qwen3-8B-Base, are trained on textual data, while the VLMs, like Qwen2.5-VL-7B-Instruct, are trained on visual data. All models are optimized using the GRPO algorithm. The results are presented in Table 3.
As shown in Table 3, all models show significant improvements on the SSMR-Bench after RL. For instance, in the Textual Modality, the Qwen3-8B-Base model initially achieved relatively low scores across all metrics, with an average of 23.18%. However, after applying GRPO to the synthetic training set, the average performance increased by 47.76%, resulting in an impressive final score of 70.9%. This marks a 13.06% improvement over the baseline Qwen3-8B-Thinking score of 57.88%. In the Visual Modality, the Qwen2.5-VL-7B-Instruct model initially recorded moderate results with an average score of 41.81%, prior to applying RLVR. Following RL training, the model’s performance improved by an average of 28.75% points, reaching a final score of 70.56%. The table also demonstrates that models with larger sizes have higher reasoning scaling potential. For example, Qwen3-8B-Base+GRPO achieves a 2% increase compared to Qwen3-4B-Base+GRPO. Additionally, while the initial performance of Qwen2.5-VL-7B-IT and Qwen2.5-VL-3B-IT is similar, after applying GRPO, Qwen2.5-VL-7B-IT outperforms Qwen2.5-VL-3B-IT by 11.56%. While our synthetic data effectively demonstrates the models’ improvements within this scenario, we also explore their applicability and potential benefits in real-world contexts.
4 From Synthetic Training to Real-World Gains
4.1 From Verifiable Music Training to Real-World Benchmarks
Table 4: Evaluation results on MusicTheoryBench (textual modality) and MMMU Music (visual modality).
Models Knowledge Reasoning Avg.
MusicTheoryBench (Textual QA)
@@ 4.1 From Verifiable Music Training to Real-World Benchmarks
Qwen3-4B-Base $21.19$ $11.22$ $16.21$
   + GRPO $55.39$ $34.69$ $45.04$
Qwen3-4B-Thinking $59.85$ $27.55$ $43.70$
@@
   + GRPO $63.20$ $36.73$ $49.97$
Qwen3-8B-Thinking $62.83$ $29.59$ $46.21$
Llama3.1-8B-IT $43.12$ $16.32$ $29.72$
   + GRPO $57.25$ $27.55$ $42.40$
MMMU Music (Visual QA, Val, Avg@8)
Qwen2.5-VL-3B-IT $18.33$ $18.33$
   + GRPO $23.33$ $23.33$
Qwen2.5-VL-7B-IT $20.00$ $20.00$
   + GRPO $26.67$ $26.67$
To assess the real-world effectiveness of our synthetic data, we conduct experiments on models trained using our synthetic data with GRPO and test them on previously established human-crafted sheet music reasoning benchmarks. These benchmarks include MusicTheoryBench (Yuan et al., 2024) and MMMU Music (Yue et al., 2024). For MusicTheoryBench, we assess models including Qwen3-4B-Base, Qwen3-8B-Base, and Llama3.1-8B-IT, comparing their performance against their original versions and the official post-trained models. Additionally, we report results for GPT-4, GPT-4 with Role Play, and GPT-4 with Chain-of-Thought (CoT) to provide further context (Yuan et al., 2024). For the MMMU Music benchmark, we test the Qwen2.5-VL-3B-IT and Qwen2.5-VL-7B-IT models, both with and without the addition of GRPO. The evaluation results are presented in Table 4.
As shown in Table 4, training on synthetic data leads to significant improvements for models on real-world datasets. For MusicTheoryBench, models like Qwen3-8B-Base and Qwen3-4B-Base demonstrate substantial gains in both knowledge and reasoning, with Qwen3-8B-Base achieving an average improvement of 21.32%, including a notable 25.28% increase in knowledge and 17.34% in reasoning. Similarly, Llama3.1-8B-IT benefits from a 14.13% increase in knowledge and 11.23% in reasoning. Besides, Qwen3-8B-Base trained with synthetic data outperforms GPT-4 and demonstrates reasoning capabilities comparable to GPT-4 with Role Play and with Chain-of-Thought (CoT). For the MMMU Music, both Qwen2.5-VL-3B-IT and Qwen2.5-VL-7B-IT show improved performance after GRPO, with increases of 5.00% and 6.67%, respectively.
4.2 Verifiable Music Training Elicits Generalizable Reasoning
Figure 5: Performance of Qwen3-8B-Base and its variants on math benchmarks and MusicTheoryBench. +GRPO (Math) is trained on a math dataset, while +GRPO (Music) is trained on our synthetic sheet music reasoning problems.
To assess the generalizability of the obtained reasoning ability beyond music, we evaluate Qwen3-8B-Base trained on synthetic sheet music data and math data across several math benchmarks (AIME24, AMC, MATH-500, Minerva, Olympiad Bench). We also include results from MusicTheoryBench for comparison with the music domain. The results are shown in Figure 5, and more details can be found in Appendix E. In the mathematical domain, training the Qwen3-8B-Base model on our synthetic dataset leads to substantial improvements. The model’s average performance increases from 33.75% to 43.79% across multiple benchmarks. However, this still falls short of the performance when trained on Math data, which achieves 52.84%. In contrast, when trained on music data, the model outperforms its math-trained counterpart on MusicTheoryBench, with a score of 49.97% compared to 33.81%.
These results confirm the effectiveness of synthesizing verifiable sheet music problems in boosting the model’s understanding and reasoning abilities within music. This reasoning ability can also generalize to math problems, although its performance is comparatively lower than that of models trained directly on mathematical data. In contrast, training on MATH data leads to notably smaller gains in sheet music reasoning. Overall, these findings suggest that automatically generating verifiable sheet music problems provides a practical and effective approach for strengthening reasoning capabilities in sheet music comprehension.
4.3 Verifiable Music Training Enhances Music Composition
@@
Table 5: Comparison of Rhythmic Consistency, with results evaluated by Gemini-2.5-Pro following the guideline in Appendix F.2.
Models RC
Sheet Music Continuation
Qwen3-8b-Base $59.00$
  + GRPO $76.00$
Qwen3-8B-Thinking $61.50$
We introduce the Rhythmic Consistency (RC), defined as the average of a binary score assigned to each generated continuation, where a sample is scored 1 if all four measures are rhythmically correct and 0 otherwise. To ensure accuracy and expertise, we collaborate with a music expert to craft the annotation guidelines for evaluating rhythmic correctness, as detailed in Appendix F.2. For evaluation, we randomly select 200 pieces of sheet music from the SSMR-Bench and generate continuations for each model. The results are then assessed by Gemini-2.5-Pro according to the established guidelines. We randomly select cases for manual verification by human experts to ensure model accuracy. Examples are provided in Appendix F.3.
@@
The results in Table 5 show that training on synthetic sheet music problems can lead to improvements in music composition. The base Qwen3-8B model achieves a score of 59.00%. After reinforcement learning on our synthetic dataset, this score dramatically increases to 76.00%, surpassing the 61.50% from the original Qwen3-8B. Although we focus only on the rhythm dimension, which is just one aspect of music, the observed improvements also highlight the potential for enhancing reasoning abilities to facilitate music composition.
5 Related Work
5.1 Sheet Music Understanding
The application of artificial intelligence to sheet music analysis has consistently been a focal point for the research community (Fornés et al., 2012; Shatri & Fazekas, 2024). Optical Music Recognition (OMR) employs AI-driven tools to interpret music notation, necessitating both specialized expertise and domain knowledge (Ma et al., 2024; Tuggener et al., 2018; Calvo-Zaragoza & Rizo, 2018). Some research in this area focuses on the recognition of handwritten music scores, introducing widely used benchmarks and datasets (Torras et al., 2024; Mayer et al., 2024), while other studies propose novel methodologies to address the underlying challenges (Rebelo et al., 2012; Rıos-Vila et al., 2024; Ríos-Vila et al., 2024; Li et al., 2023).
While significant progress has been made, some researchers are expanding the study to Sheet Music Question Answering (QA), moving beyond traditional music recognition and image modalities. MMMU (Yue et al., 2024) provides music QA, highlighting the challenges of sheet music QA based on images. Meanwhile, there is growing interest in using LLMs for symbolic music (Wu et al., 2024; Wang et al., 2025; Qu et al., 2025), where sheet music is represented in textual formats, such as the widely used ABC notation.  Yuan et al. (2024) introduces a MusicTheoryBench that includes knowledge and reasoning categories to evaluate the music understanding abilities of LLMs. However, the manually crafted nature of these benchmarks hinders data scalability.
5.2 Reinforcement Learning Through Synthetic Data
@@ 6 Conclusion
This work introduces the first framework to programmatically generate verifiable sheet music problems using music theory rules, providing both textual and visual modalities. Based on this framework, we introduce SSMR-Bench and a corresponding training set in both modalities. Experimental results demonstrate that the synthetic data effectively enhances the reasoning capabilities of LLMs and MLLMs in sheet music tasks, underscoring its potential as a valuable resource. Moreover, the results also highlight the promise of this enhanced ability to facilitate music composition. Our work demonstrates the potential of programmatically generating verifiable sheet music problems to evaluate and improve the abilities of LLMs and MLLMs in sheet music, while also highlighting promising avenues for future research toward AI musicianship.
ETHICS STATEMENT
We believe there are no major ethical concerns associated with our work. The study is exclusively focused on the music domain, specifically sheet music reasoning, and does not involve any applications that could be deemed sensitive or harmful.
Reproducibility statement
We make our code and the generated data available at Anonymous GitHub . This repository includes our framework for synthesizing sheet music reasoning problems, along with the corresponding evaluation scripts. Besides, we provide the generated data, covering both textual and visual modalities.
@@ Appendix A Dataset Details
  • Rhythm Questions assess a model’s ability to understand sheet music rhythms, requiring accurate calculation of note durations and recognition of overall temporal patterns.
@@
A.1 Question Class Descriptions
@@ A.1.1 Domain: Scale
A.1.2 Domain: Rhythm
@@
[…]
Ground_truth:B
A.1.3 Domain: Interval
@@ Appendix B Training Template
Appendix C Results Analysis by Question Type
Figure 6: Performance comparison of various selected LLMs and MLLMs on different music-related question types. The upper plot shows the performance of LLMs evaluated on the textual modality, while the lower plot displays the performance of MLLMs in the visual modality. The performance is assessed across nine distinct question categories, such as Bar Placement, Chord Identification (Chord ID), and Scale Identification (Scale ID). Detailed descriptions of these question types can be found in Appendix 6.
The upper plot of Figure 6 illustrates the performance of LLMs in the textual modality. After reinforcement learning on 8,000 synthetic music reasoning samples through GRPO, Qwen3-8B-Base outperforms Qwen3-8B in 2/3 of the question types, despite the latter undergoing extensive data training. Additionally, Qwen3-8B-Base+GRPO achieves comparable performance to Deepseek-V3 and Qwen3-32B in aspects such as "Time Sig", "Interval No", and "Note Comp".
Appendix D Training Dynamics of RLVR on Synthetic Sheet Music Reasoning Data
Appendix E Detailed Results on Math Benchmarks for RL with Synthetic Sheet Music Data
Models Math Benchmarks
AIME24 AMC MATH-500 Minerva Olympiad Bench
Qwen3-8B-Thinking $50.41$ $71.08$ $91.60$ $45.96$ $60.89$
Qwen3-8B-Base $8.75$ $39.16$ $66.00$ $25.37$ $30.22$
+ GRPO (Math) 25.8325.83(+17.08) 62.0462.04(+22.88) 85.4085.40(+19.40) 40.4440.44(+15.07) 50.5250.52(+20.30)
+ GRPO (Music) 13.7513.75(+5.00) 52.1152.11(+12.95) 79.4079.40(+13.40) 34.1934.19(+8.82) 40.4440.44(+10.22)
Appendix F Sheet Music Generation and Evaluation
F.1 Prompt for sheet music generation
[…]
Figure 8: Prompt for Sheet Music Continuation
F.2 The Guideline for Rhythm Consistency Evaluation
The guideline shown in Figure 9 and Figure 10 outlines a systematic approach to assess the rhythmic accuracy and consistency of automatically generated sheet music when using ABC 2.1 notation. The evaluation process is divided into two primary stages: a foundational syntax check and a detailed rhythmic integrity analysis.
[…]
F.3 Case Evaluation Results with Expert Judgement
To ensure the correctness of Gemini-2.5-Pro’s judgments, we present several cases of the model’s evaluation results alongside expert verification. For each case, the prompt, the model-generated sheet music, Gemini-2.5-Pro’s verification process, and the human expert’s judgment are provided.
Figure 11: Sheet Music Continuation—Case 1. Model: Qwen3-8B-Base. Includes the prompt, model response, Gemini-2.5-Pro verification, and human evaluation result.
Figure 12: Sheet Music Continuation—Case 2. Model: Qwen3-8B-Base. Includes the prompt, model response, Gemini-2.5-Pro verification, and human evaluation result.
Figure 13: Sheet Music Continuation—Case 3. Model: Qwen3-8B-Base. Includes the prompt, model response, Gemini-2.5-Pro verification, and human evaluation result.
Figure 14: Sheet Music Continuation—Case 4. Model: Qwen3-8B-Base+GRPO. Includes the prompt, model response, Gemini-2.5-Pro verification, and human evaluation result.
Figure 15: Sheet Music Continuation—Case 5. Model: Qwen3-8B-Base+GRPO. Includes the prompt, model response, Gemini-2.5-Pro verification, and human evaluation result.
Figure 16: Sheet Music Continuation—Case 6. Model: Qwen3-8B-Base+GRPO. Includes the prompt, model response, Gemini-2.5-Pro verification, and human evaluation result.
Figure 17: Sheet Music Continuation—Case 7. Model: Qwen3-8B-Thinking. Includes the prompt, model response (excluding content within <think></think>), Gemini-2.5-Pro verification, and human evaluation result.
Figure 18: Sheet Music Continuation—Case 8. Model: Qwen3-8B-Thinking. Includes the prompt, model response (excluding content within <think></think>), Gemini-2.5-Pro verification, and human evaluation result.
Figure 19: Sheet Music Continuation—Case 9. Model: Qwen3-8B-Thinking. Includes the prompt, model response (excluding content within <think></think>), Gemini-2.5-Pro verification, and human evaluation result.